# Cross-Calibration and First Vertical ECE Measurement of Electron Energy Distribution in the TCV Tokamak

### v1

@@ 4.2 Runaway electron measurement
Figure 16: Time histories of runaway electron discharge #$\#73002$ featuring typical plasma parameters plus measurements of X-mode radiation from the $78-96$ GHz radiometer.
@@ 5.2 Inferring the electron energy distribution from V-ECE measurements
The ratios of X- to O-mode intensities measured during the heating phase of the ECCD experiment #$\#72644$ are shown in Figure 18. The O-mode intensities were measured with the same radiometer, in the frequency range $96-114$ GHz, with a twin discharge while the wire grid polariser was rotated 90. The energies are estimated, assuming harmonic overlap can be neglected, and that non-thermal contribution to the radiation comes essentially from the third harmonic emission.
@@
For the lower energies, The ratio decreases with increasing $\theta_{L}$, suggesting that the parallel energy is been preferentially enhanced after the first launcher steps. This could mean that the interaction between the lower energy electrons and the ECH wave, which in principle enhances perpendicular energy, decreases with increasing $N_{\parallel}$. In that case, the low energy electrons would have their energy enhanced from collisions with the higher energy electrons. The enhancement of the parallel energy through collisions may be responsible for the reduction of X- to O-mode power ratio. At the lower energies, the ratio is higher for smaller $N_{\parallel}$ and launcher angle $\theta_{L}$. This trend is inverted at the higher energies, where the polarisation ratios have lower values in general. It is observed that small $N_{\parallel}$ leads to stronger enhancement in the perpendicular direction, explaining the higher ratio of the X- to the O- mode intensities. At higher $N_{\parallel}$ (approximately between $1.4$ s and $1.6$ s), the higher energies are more enhanced in the perpendicular direction, consistently with the expectations.
@@
The ratio of the intensities of the X- to the O-mode intensity for the runaway electron case #$\#73000$ is shown in Figure 19. The ratio sharply decreases when the runaway beam starts, confirming that it is the parallel energy of the electrons that is being enhanced, in particular due to the toroidal electric field acceleration. In general the ratio is lower for runaway electron radiation, at each measured frequency, compared to the suprathermal electrons generated by ECCD in the case presented in Figure 18. Note that the polarisation ratios are obtained from the raw data and are free from the uncertainties in the calibration.
@@ 6 Conclusion
The work described in this article can inspire future research on V-ECE diagnostics and ECE instrument calibration, especially for upcoming fusion machines like ITER, SPARC, and DEMO. Since venting of these next-generation devices will be limited, techniques that allow regular calibration of ECE diagnostics without machine venting, like the one demonstrated in this study, will be beneficial. The calibration technique was shown to be applicable to arbitrary lines of sight and optical depths of the plasma. Its accuracy on TCV has relied on scenarios almost free of multiply reflected radiation, regardless of the presence of a viewing dump. This V-ECE result, reproducible in other machines, has been demonstrated for the first time in this work. The results presented in this article can contribute to reviving interest in non-thermal plasma diagnostics using ECE, which will be beneficial for Tokamak operations in current and future fusion devices.

### v2

@@ 4.2 Runaway electron measurement
Figure 16: Time histories of V-ECE measurement in runaway electron discharge #$\#73002$ featuring typical plasma parameters plus measurements of X-mode radiation from the $78-96$ GHz radiometer.
@@ 5.2 Inferring the electron energy distribution from V-ECE measurements
The ratios of X- to O-mode intensities measured during the heating phase of the ECCD experiment #$\#72644$ are shown in Figure 18. The O-mode intensities were measured using the same radiometer, in the frequency range 96–114 GHz, with a twin discharge (an exact repeat of discharge #$\#72644$), while the wire grid polariser was rotated 90° in the plane parallel to the grid. The energies are estimated, assuming harmonic overlap can be neglected, and that non-thermal contribution to the radiation comes essentially from the third harmonic emission.
@@
For the lower energies, The ratio decreases with increasing $\theta_{L}$, suggesting that the parallel energy is been preferentially enhanced after the first launcher steps. This could mean that the interaction between the lower energy electrons and the ECH wave, which in principle enhances perpendicular energy, decreases with increasing $\left|N_{\parallel}\right|$. In that case, the low energy electrons would have their energy enhanced from collisions with the higher energy electrons. The enhancement of the parallel energy through collisions may be responsible for the reduction of X- to O-mode power ratio. At the lower energies, the ratio is higher for smaller $\left|N_{\parallel}\right|$ and launcher angle $\theta_{L}$. This trend is inverted at the higher energies, where the polarisation ratios have lower values in general. It is observed that small $\left|N_{\parallel}\right|$ leads to stronger enhancement in the perpendicular direction, explaining the higher ratio of the X- to the O- mode intensities. At higher $\left|N_{\parallel}\right|$ (approximately between $1.4$ s and $1.6$ s), the higher energies are more enhanced in the perpendicular direction, consistently with the expectations.
@@
The ratio of the intensities of the X- to the O-mode intensity for the runaway electron case #$\#73000$ is shown in Figure 19. The ratio sharply decreases when the runaway beam starts, confirming that it is the parallel energy of the electrons that is been enhanced, in particular due to the toroidal electric field acceleration. In general the ratio is lower for runaway electron radiation, at each measured frequency, compared to the suprathermal electrons generated by ECCD in the case presented in Figure 18. Note that the discussed polarisation ratios are free from the calibration uncertainties mentioned in Section 3.6. The ratios are obtained from the radiometers’ raw data, accounting for the radiometers’ insertion losses. The insertion losses were determined by rotating the wire grid polariser by 45 with the plasma as a source, and by using an external power source from a Vector Network Analyzer.
@@ 6 Conclusion
The work described in this article displays important advancements over previously published results and has the potential to inspire future research on V-ECE diagnostics and calibration of ECE instruments. A key strength of this work is the interpretation of diagnostic results almost free from thermal background radiation. This achievement is made possible by better control over multiple wall reflections and represents an innovation compared to previous studies, where viewing dumps were considered sine qua non conditions for operating the diagnostic. The control of background radiation, demonstrated for the first time in this work has notably enabled a more accurate cross-calibration of V-ECE with Thomson Scattering. Techniques that allow regular calibration of ECE diagnostics without requiring machine venting will be beneficial for next-generation machines such as ITER, SPARC, or DEMO. More broadly, the results presented in this article could help revive interest in fast electron diagnostics in Tokamaks using ECE. As demonstrated on TCV, V-ECE can play a key role in the detection and mitigation of low-energy runaway electrons, bringing us a step closer to achieving fusion energy.